OpenStudy (anonymous):

What is the solution set of |x – 4| + 7 = 4?

OpenStudy (anonymous):

subtract 7 from both side

OpenStudy (anonymous):

subtract the 7 from both sides, then solve it the same way you did the last one.

OpenStudy (anonymous):

I don't think the equation has any real solutions.

OpenStudy (anonymous):

|x-4|+7=4 Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides. |x-4|=-7+4 Add 4 to -7 to get -3. |x-4|=-3 Remove the absolute value term. This creates a \ on the right-hand side of the equation because |x|=\x. x-4=\(-3) Set up the portion of the \ solution. x-4=-3 Move all terms not containing x to the right-hand side of the equation. x=1 Set up the - portion of the \ solution. x-4=-(-3) Multiply -1 by the -3 inside the parentheses. x-4=3 Since -4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. x=4+3 Add 3 to 4 to get 7. x=7 The solution to the equation includes both the positive and negative portions of the solution. x=1,7

OpenStudy (anonymous):

x = 1, 7

OpenStudy (anonymous):

no chance

OpenStudy (anonymous):

That's true actually.

OpenStudy (anonymous):

yup,no solutions

OpenStudy (anonymous):

kushash.. your script doesn't work for this.

OpenStudy (anonymous):

why polpak .

OpenStudy (anonymous):

Because |x - 4| can not equal -3. Ever.

OpenStudy (anonymous):

where i did wrong tell me

OpenStudy (anonymous):

I did..

OpenStudy (anonymous):

what u mean , given |x-4|+7=4 then can't we transpose +7 to right hand side getting |x-4|=4-7 |x-4|=-3